# Hyperspectral Image Data Reduction for Endmember Extraction

### v1

@@ 5.2 Performance Evaluation of DRS
First, we evaluated the performance of the DRS module incorporated into REDIC using synthetic datasets. The datasets were derived from the three real HSI datasets, as described in Section 5.1, following the procedure presented in Section VII-B of [Miz25]. To make this paper self-contained, we summarize the procedure here. The endmember signatures for each HSI dataset were identified in [Zhu17], and we used them to generate the synthetic datasets. Let $A^{\mathrm{real}}\in\mathbb{R}^{d\times n}$ denote the HSI matrix for a given HSI dataset, and let $\bm{w}_{1}^{\mathrm{ident}},\ldots,\bm{w}_{r}^{\mathrm{ident}}$ represent the identified endmember signatures. The following procedure was used to generate $W\in\mathbb{R}^{d\times r}$, $H\in\mathbb{R}^{r\times n}$, and $V\in\mathbb{R}^{d\times n}$:
@@ 5.3 Performance Evaluation of REDIC
Next, we evaluated the performance of REDIC using the three real HSI datasets described in Section 5.1. The following metric was employed to assess the accuracy of endmember extraction. Let $A=[\bm{a}_{1},\ldots,\bm{a}_{n}]\in\mathbb{R}^{d\times n}$ denote the HSI matrix for a given dataset, and let $\bm{w}_{1}^{\mathrm{ident}},\ldots,\bm{w}_{r}^{\mathrm{ident}}\in\mathbb{R}^{d}$ represent the endmember signatures identified in [Zhu17]. We compute the column indices $i_{1},\ldots,i_{r}$ of $A$ corresponding to $\bm{w}_{1}^{\mathrm{ident}},\ldots,\bm{w}_{r}^{\mathrm{ident}}$ under the MRSA metric as
@@
This experiment was based on the EEHT-C experiment on the Urban dataset presented in [Miz25], where a data-specific preprocessing technique was applied prior to executing EEHT-C to enhance its performance (see Section VII-C of [Miz25] for details). The preprocessing technique involves two parameters, $\phi$ and $\omega$. In this experiment, before executing endmember extraction, we applied the same preprocessing to the Urban dataset, using $(\phi,\omega)=(0.4,0.1)$ as in [Miz25]. This procedure eliminated 2995 columns from the original HSI matrix, which contained 94249 columns, resulting in a total of 91254 columns. For Jasper Ridge and Samson, no such preprocessing was applied.
@@
We then present the experimental results for REDIC. The REDIC and LP methods were applied to each HSI dataset, and their MRSA scores and computational times were compared. The LP method examined here is identical to EEHT-C proposed in [Miz25], whose procedure was reviewed in Section 4. The parameters $\lambda$ and $\tau$ in REDIC were set as follows: $\tau$ took values in $\{1,5\}$, and $\lambda$ was varied in the same manner as in the first experiment described in Section 5.2. When $\lambda>0$, REDIC was executed 50 times for each of the three HSI datasets, and the mean and standard deviation of the MRSA scores and elapsed times across the 50 runs were calculated. Table 3 summarizes the experimental results. We can make the following observations from these results:
@@
The previous version of the MATLAB code for the LP method (i.e., EEHT-C) developed in [Miz25] required 10 hours to process the Urban dataset. In contrast, the reorganized and speed-improved version of the code used in this experiment completed the same task in 8 hours.
@@ 7 Concluding Remarks
We conclude this paper by suggesting directions for future research. As observed in Section 5.3, there are noticeable gaps between the MRSA distances of the outputs produced by DRS and the MRSA scores obtained by REDIC with $\lambda=0$, indicating that the extraction accuracy of REDIC still leaves room for improvement. One possible approach to addressing this limitation is to use a combinatorial method in place of the LP method within REDIC. Gillis [Gil19] studied a combinatorial approach for a slightly generalized separable NMF problem and showed theoretically that it exhibits strong robustness to noise. This suggests that combinatorial methods may achieve superior endmember extraction accuracy. Due to their high computational cost, applying such methods directly to endmember extraction is impractical. However, by integrating them with DRS, it may be possible to achieve higher accuracy while maintaining reasonable computational time, potentially outperforming the current version of REDIC.

### v2

@@ 5.2 Performance Evaluation of DRS
First, we evaluated the performance of the DRS module incorporated into REDIC using synthetic datasets. The datasets were derived from the three real HSI datasets, as described in Section 5.1, following the procedure presented in Section VII-B of [25]. To make this paper self-contained, we summarize the procedure here. The endmember signatures for each HSI dataset were identified in [29], and we used them to generate the synthetic datasets. Let $A^{\mathrm{real}}\in\mathbb{R}^{d\times n}$ denote the HSI matrix for a given HSI dataset, and let $\bm{w}_{1}^{\mathrm{ident}},\ldots,\bm{w}_{r}^{\mathrm{ident}}$ represent the identified endmember signatures. The following procedure was used to generate $W\in\mathbb{R}^{d\times r}$, $H\in\mathbb{R}^{r\times n}$, and $V\in\mathbb{R}^{d\times n}$:
@@ 5.3 Performance Evaluation of REDIC
Next, we evaluated the performance of REDIC using the three real HSI datasets described in Section 5.1. The following metric was employed to assess the accuracy of endmember extraction. Let $A=[\bm{a}_{1},\ldots,\bm{a}_{n}]\in\mathbb{R}^{d\times n}$ denote the HSI matrix for a given dataset, and let $\bm{w}_{1}^{\mathrm{ident}},\ldots,\bm{w}_{r}^{\mathrm{ident}}\in\mathbb{R}^{d}$ represent the endmember signatures identified in [29]. We compute the column indices $i_{1},\ldots,i_{r}$ of $A$ corresponding to $\bm{w}_{1}^{\mathrm{ident}},\ldots,\bm{w}_{r}^{\mathrm{ident}}$ under the MRSA metric as
@@
This experiment was based on the EEHT-C experiment on the Urban dataset presented in [25], where a data-specific preprocessing technique was applied prior to executing EEHT-C to enhance its performance (see Section VII-C of [25] for details). The preprocessing technique involves two parameters, $\phi$ and $\omega$. In this experiment, before executing endmember extraction, we applied the same preprocessing to the Urban dataset, using $(\phi,\omega)=(0.4,0.1)$ as in [25]. This procedure eliminated 2995 columns from the original HSI matrix, which contained 94249 columns, resulting in a total of 91254 columns. For Jasper Ridge and Samson, no such preprocessing was applied.
@@
We then present the experimental results for REDIC. The REDIC and LP methods were applied to each HSI dataset, and their MRSA scores and computational times were compared. The LP method examined here is identical to EEHT-C proposed in [25], whose procedure was reviewed in Section 4. The parameters $\lambda$ and $\tau$ in REDIC were set as follows: $\tau$ took values in $\{1,5\}$, and $\lambda$ was varied in the same manner as in the first experiment described in Section 5.2. When $\lambda>0$, REDIC was executed 50 times for each of the three HSI datasets, and the mean and standard deviation of the MRSA scores and elapsed times across the 50 runs were calculated. Table 3 summarizes the experimental results. We can make the following observations from these results:
@@
The previous version of the MATLAB code for the LP method (i.e., EEHT-C) developed in [25] required 10 hours to process the Urban dataset. In contrast, the reorganized and speed-improved version of the code used in this experiment completed the same task in 8 hours.
@@ 7 Concluding Remarks
We conclude this paper by suggesting directions for future research. As observed in Section 5.3, there are noticeable gaps between the MRSA distances of the outputs produced by DRS and the MRSA scores obtained by REDIC with $\lambda=0$, indicating that the extraction accuracy of REDIC still leaves room for improvement. One possible approach to addressing this limitation is to use a combinatorial method in place of the LP method within REDIC. Gillis [18] studied a combinatorial approach for a slightly generalized separable NMF problem and showed theoretically that it exhibits strong robustness to noise. This suggests that combinatorial methods may achieve superior endmember extraction accuracy. Due to their high computational cost, applying such methods directly to endmember extraction is impractical. However, by integrating them with DRS, it may be possible to achieve higher accuracy while maintaining reasonable computational time, potentially outperforming the current version of REDIC.